How much would you pay for a perpetual bond that pays an annual coupon of $100 per year and yields on competing instruments are 10%? You would pay $(Round your response to the nearest penny.) If competing yields are expected to change to 8%, what is the current yield on this same bond assuming that you paid $1,000? The current yield is %. (Round your response to the nearest integer) If you sell this bond in exactly one year, having paid $1,000, and received exactly one coupon payment, what is your total return if competing yields are 8%? Your total return is %. (Round your response to two decimal places)
Answer #1

Coupon amount=C=$100

Expected return=r=10%

Fair price of bond, P=Present value of perpetual coupon payments

P=C/r=100/10% =$1000

If competing yield changes to 8%, price of bond is

P=C/r=100/8% =$1250

Current yield=C/current price=100/1250=8%

Current price=S=$1250

Purchase price=P=$1000

Coupon payment received=C=$100

Holding period return=[(S-P)+C]/P=[(1250-1000)+100]/1000=35.00%